### v1

# ArchesClimate: Probabilistic Decadal Ensemble Generation With Flow Matching

Abstract
Climate projections have uncertainties related to components of the climate system and their interactions. A typical approach to quantifying these uncertainties is to use climate models to create ensembles of repeated simulations under different initial conditions. Due to the complexity of these simulations, generating such ensembles of projections is computationally expensive. In this work, we present ArchesClimate, a deep learning-based climate model emulator that aims to reduce this cost. ArchesClimate is trained on decadal hindcasts of the IPSL-CM6A-LR climate model at a spatial resolution of approximately 2.5x1.25 degrees. We train a flow matching model following ArchesWeatherGen \@BBOPcitep\@BAP\@BBN(Couairon et al., 2024)\@BBCP, which we adapt to predict near-term climate. Once trained, the model generates states at a one-month lead time and can be used to auto-regressively emulate climate model simulations of any length. We show that for up to 10 years, these generations are stable and physically consistent. We also show that for several important climate variables, ArchesClimate generates simulations that are interchangeable with the IPSL model. This work suggests that climate model emulators could significantly reduce the cost of climate model simulations.
[…]
Journal of Advances in Modeling Earth Systems (JAMES)
ARCHES, INRIA, Paris, France UMR LOCEAN, IPSL, Sorbonne Université, IRD, CNRS, MNHN, Paris, France Department of Computer Science, University of Colorado Boulder, USA
[…]
Graham Clynegraham.clyne@inria.fr
{keypoints}
We present ArchesClimate, an AI model for probabilistically emulating a climate model at a monthly timescale
ArchesClimate simultaneously generates both oceanic and atmospheric dynamics that can be autoregressively generated for 10 years
We propose a set of evaluations of both the statistics and the physical properties of the AI-generated ensembles
Plain Language Summary
Climate modeling enables us to understand the impacts of a changing climate. Simulations from climate models can address a wide range of questions, including near-term (one to ten years) climate evolution to allow for informed and immediate policy decisions. To increase confidence in predictions from climate models, many simulations are integrated to form a probabilistic picture of the climate. These simulations require significant computational resources because of the complexity of the climate models used. Here we present ArchesClimate, a model to replicate the behavior of a climate model. Using machine learning, we are able to significantly reduce the computational cost of climate models. ArchesClimate generates both atmospheric and oceanic dynamics at a monthly temporal resolution and approximately 2.5x1.25 degree spatial resolution for up to 10 years.
1 Introduction
Ensemble generation is an important tool to investigate the climate at any timescale. Given the chaotic nature of the atmospheric and ocean dynamics, small variations in initial conditions can evolve to vastly different states. As such, ensembles of climate model simulations can be seen as distributions of possible future climates. By repeatedly sampling the simulated climate, we enable probabilistic analyses that support the investigation of a broad range of problems, including extreme event attribution and uncertainty quantification \@BBOPcitep\@BAP\@BBN(Fyfe et al., 2017; Fischer et al., 2023)\@BBCP.
Furthermore, generating ensembles can significantly contribute to separating the internal variability from the effect of other sources of uncertainty, such as those associated with external forcings \@BBOPcitep\@BAP\@BBN(Maher et al., 2021; Eade et al., 2014)\@BBCP. Internal variability in climate models refers to the natural fluctuations in the climate system that occur without variations from external forcings (e.g. greenhouse gases or solar irradiation). These fluctuations arise from complex interactions between the atmosphere, ocean, land, and ice (e.g. El Niño or the North Atlantic Oscillation) — and can cause variations in climate at all time scales. Understanding internal variability improves confidence in detecting and projecting human-caused climate change.
Traditionally, ensemble members are generated by perturbing a set of initial conditions and running several instances of a climate model \@BBOPcitep\@BAP\@BBN(Deser et al., 2024)\@BBCP. This process is computationally expensive and thereby limits the widespread adoption of large ensemble generation \@BBOPcitep\@BAP\@BBN(Smith et al., 2019)\@BBCP. To reduce the cost of generating such ensembles, many atmospheric emulators have been made with increasing success. In \@BBOPcite\@BAP\@BBNBrenowitz et al. (2025)\@BBCP, the authors introduce a diffusion-based machine learning model trained on multiple atmospheric datasets to learn both weather and climate dynamics. While their approach demonstrates potential in generating realistic instantaneous states, it does not produce temporally consistent climatic sequences since the model is not auto-regressive, a seemingly critical requirement for long-term climate modeling. In contrast, \@BBOPcite\@BAP\@BBNWatt-Meyer et al. (2024)\@BBCP develops the ACE2 model, a deterministic emulator of ERA5 which can stably generate sequential states for up to 1000 years, emulating atmospheric dynamics that respond to forcings. While ACE2 marks a breakthrough in long-term atmospheric emulation, it is limited by its focus on the atmosphere alone, leaving coupled ocean–atmosphere processes and variability unresolved. Moreover, its deterministic design may underrepresent internal variability and uncertainty. Generalization beyond the present climate regime, particularly under future or paleoclimate conditions, also remains uncertain.
Several efforts have also been made to emulate oceanic components of climate models. \@BBOPcite\@BAP\@BBNGuo et al. (2024)\@BBCP proposes a global ocean emulator trained on climate simulations that produces stable decadal simulations. Complementing this approach, \@BBOPcite\@BAP\@BBNDheeshjith et al. (2025)\@BBCP introduces Samudra, a deterministic model that focuses on learning spatiotemporal coherence in ocean circulation patterns. While it performs well at depth and captures inter-annual variability, it does not respond to a changing climate.
Each of these approaches incorporates design decisions tailored to their specific objectives, but our goal is distinct: to efficiently emulate the IPSL-CM6A-LR at decadal timescales using probabilistic modeling across both oceanic and atmospheric domains. We introduce ArchesClimate, an AI-driven probabilistic emulator that aims to learn climatic processes at a monthly temporal resolution. We adapted ArchesWeatherGen and trained it to emulate the IPSL-CM6A-LR model at approximately 2.5x1.25 degree resolution. We build on ArchesWeatherGen \@BBOPcitep\@BAP\@BBN(Couairon et al., 2024)\@BBCP, a state-of-the-art AI-based numerical weather prediction model, which is based on PanguWeather \@BBOPcitep\@BAP\@BBN(Bi et al., 2022)\@BBCP. ArchesWeatherGen provides a computationally efficient solution for ensemble generation that is adapted to climatic timescales.
We use a dataset from the Decadal Climate Prediction Project (DCPP), a Model-Intercomparison Project aimed at improving predictability and understanding climate at the decadal timescale \@BBOPcitep\@BAP\@BBN(Boer et al., 2016)\@BBCP. It comprises ensembles of 10 members with a duration of 10 years, starting every year between 1960 and 2015. We use flow matching as our training scheme, a recent generative technique that learns a function to map one distribution to another distribution, usually from a multivariate normal distribution to the target data distribution \@BBOPcitep\@BAP\@BBN(Lipman et al., 2023)\@BBCP.
We distinguish between an emulator of the input/output functionality of the climate model versus an emulator of how a climate model evolves in climate state. The former describes a machine learning model that takes in boundary conditions or forcings and outputs full climate states, replacing the functionality of a climate model. The latter describes a machine learning model that learns from the output of a climate model and learns dynamics from this output, often augmenting the emulated climate model. In this research, we do the latter, and therefore our setting uses data from the IPSL-CM6A-LR for both initial conditions and training data. If we worked directly from the forcings and did not use an initial state generated from IPSL-6CMA-LR, we could instead attempt to replace the functioning of IPSL-6CMA-LR.
The manuscript is structured as follows: Section 2.1 and Section 2.2 present the dataset and architecture used in the research, Section 2.3 and Section 2.4 outline how we train and generate data, and Section 3 explains the experiments and their results. Section 4 discusses conclusions and next steps.
2 Materials and Methods
2.1 Dataset
We use generated states from the IPSL-CM6A-LR coupled climate model \@BBOPcitep\@BAP\@BBN(Boucher et al., 2020)\@BBCP submission to the DCPP \@BBOPcitep\@BAP\@BBN(Boer et al., 2016)\@BBCP. DCPP aims at exploring decadal climate prediction, its predictability and variability. We use data from hindcastA, an experiment that performs 10-year hindcasts starting from 1960 to the present. Ensembles are initialized every year on January 1st from 1960-2015. For example, the dataset contains a 10-member ensemble from 1960-1970 and another 10-member ensemble for 1961-1971. There is therefore significant temporal overlap between each ensemble. We use monthly outputs that are averaged values, corresponding to a dataset of approximately 70,000 simulated months. The 10 members have initial conditions generated by an assimilated run using observed sea surface temperature and sea surface salinity nudging over this period \@BBOPcitep\@BAP\@BBN(Estella-Perez et al., 2020; Servonnat et al., 2015; Deser et al., 2024)\@BBCP. This perturbation is enough to generate variability in the ensemble while keeping some features leading to decadal predictability unchanged. We will refer to the dataset as IPSL-DCPP hereafter.
We select a subset of the available outputs of the IPSL-CM6A-LR, omitting a large portion of the available output. We choose not to account for the vertical structure of the ocean to reduce the size of the input state and instead use the ocean heat content that represents the vertically integrated heat at different depths. We use 10 surface variables, 7 oceanic variables and 7 atmospheric variables with 4 pressure levels (250, 500, 750, 800 hPa). As we are constrained computationally, we choose instead to use this subset of variables as a proof of concept to capture different climatic states at a monthly timescale.
The variables laid out in Table 1 capture important atmospheric and oceanic dynamics, including the heat and water flux between the two domains. The variables net_flux (total heat exchange between atmosphere and ocean, see A for a detailed description) and evspsbl (evaporation) represent the interactions between the two domains. The resolution of the atmospheric model is 144x143 (lon x lat, roughly 2.5x1.25 degrees), and the resolution of the ocean model is 362x332 (lon x lat, roughly 1 degree). The oceanic variables are re-gridded onto the regular atmospheric grid using a nearest-neighbour interpolation. For oceanic variables, we mask grid points over land with zeros. We also use the atmospheric concentration of four greenhouse gases (CO2, CH4, CFC11eq, N2O) and solar irradiance. from input4mips, a data repository of standardized boundary conditions used in model-intercomparison projects \@BBOPcitep\@BAP\@BBN(Meinshausen & Nicholls, 2018; Lurton et al., 2020)\@BBCP.
Figure 1: On the right, a visualization of one state ($X_{t}$) from ISPL-DCPP, with surface and oceanic variables separated from atmospheric variables. External forcings are shown as a vector to the right.
Table 1: Variables in IPSL-DCPP
Variable Name Long Name
Surface Variables
clt Total cloud cover percentage
hurs Near-surface relative humidity
huss Near-surface specific humidity
pr Precipitation
ps Surface air pressure
tasmax Daily maximum near-surface air temperature
tasmin Daily minimum near-surface air temperature
tas Near-surface air temperature
evspsbl Evaporation including sublimation and transpiration
sfcWind Near-surface wind speed
net_flux Total positive downward flux between ocean and atmosphere.
psl Sea level pressure
Ocean Variables
thetaot2000 Vertically-averaged potential temperature at 0-2000m
thetaot700 Vertically-averaged potential temperature at 0-700m
thetaot300 Vertically-averaged potential temperature at 0-300m
t20d Depth of 20 degree celsius isotherm
tos Sea surface temperature
Atmospheric Variables at 250, 500, 700, 850 hPa
hur Relative humidity
hus Specific humidity
ta Air temperature
ua Eastward wind
va Northward wind
wap Omega
zg Geopotential height
Forcings
CO2 Atmospheric carbon dioxide (Yearly, Non-spatial)
CH4 Atmospheric methane (Yearly, Non-spatial)
N2O Nitrous oxide (Yearly, Non-spatial)
CFC12eq Trichlorofluoromethane (Yearly, Non-spatial)
SSI Spectral solar irradiance (Daily, Non-Spatial)
2.2 Architecture
Here we describe the architecture of ArchesClimate. The backbone is identical to ArchesWeatherGen \@BBOPcitep\@BAP\@BBN(Couairon et al., 2024)\@BBCP. ArchesWeatherGen is a machine learning model based on PanguWeather that uses a hierarchical vision transformer and uses shifted-window attention for efficient and scalable modeling of visual features \@BBOPcitep\@BAP\@BBN(Bi et al., 2022; Liu et al., 2021; Vaswani et al., 2017)\@BBCP. ArchesWeatherGen uses an ensemble of deterministic models to predict the ensemble mean at the following timestep (6 hours) and generates a probabilistic ensemble of forecasts of the residual to predict the residual between the mean state and true state at the following timestep.
To extend the ArchesWeatherGen approach to the decadal climate prediction domain, we make the following changes:
  • We include greenhouse gas and solar irradiance forcings using conditional layer normalization as done in \@BBOPcite\@BAP\@BBNChen et al. (2021)\@BBCP. The forcings listed in Table 1 are included as parameters.
  • We remove axial attention to facilitate modeling of multiple domains. Axial attention breaks down attention into multiple 1D attentions along different axes (e.g., height and width separately for images). At the cost of increased complexity, we leave it to the model to determine the relationship between each domain and its interactions \@BBOPcitep\@BAP\@BBN(Ho et al., 2019)\@BBCP.
  • We increase the embedded dimension to 4 times the size of ArchesWeatherGen to account for this increased model complexity.
  • We remove the skip connection to accommodate this large embedded dimension. While skip connections would preserve features and expressive capacity in the model, the total size of the model would exceed our computational resources.
  • We omit per-variable weighting in the loss function while keeping latitude weighting. Previous studies of weather emulators have included per-variable weighting \@BBOPcitep\@BAP\@BBN(Bi et al., 2022)\@BBCP, but given the different nature of our variable set, which contains ocean and atmospheric variables at a monthly timescale, we cannot reuse those weightings.
There are also changes to the training of ArchesClimate, which can be found in Section 2.3.
2.2.1 Forcings Included In ArchesClimate
To incorporate forcings in ArchesClimate, we adopt conditional layer normalization following \@BBOPcitep\@BAP\@BBN(Chen et al., 2021)\@BBCP. In this approach, each scalar forcing value is first passed through an embedding layer, which produces parameters that rescale and shift the outputs of a standard layer normalization. The forcings are CO2, CH4, CFC11eq, N2O and SSI as described in Table 1. We apply conditional layer normalization in all transformer blocks of ArchesClimate, enabling the model to integrate conditioning signals at both global and local levels. By introducing greenhouse gas forcings through this mechanism, the model can learn flexible relationships that adapt dynamically to different forcing scenarios.
2.3 Training
Following ArchesWeatherGen, we train both a deterministic and generative model to predict the next state at timestep $t+\delta$, where t is a timestep in IPSL-DCPP and $\delta$ is one month. A deterministic model, $f_{\theta}$, is trained to predict $X_{t+\delta}$ from $X_{t}$ where $X$ is a climatic state of IPSL-DCPP (see Figure 1). In this climate state, we include the forcings listed in Table 1 shown as $\text{Forcings}_{t}$ in Figure 2. They are included in both $f_{\theta}$ and $g_{\theta}$ via conditional layer normalization \@BBOPcitep\@BAP\@BBN(Chen et al., 2021)\@BBCP. We then train a generative model $g_{\theta}$ to predict the residual of the next state
$r_{t+\delta}=\frac{X_{t+\delta}-f_{\theta}(X_{t})}{\sigma}$ (1)
where $\sigma$ is the standard deviation of the residual ($X_{t+\delta}-f_{\theta}(X_{t})$) of the training dataset.
We use flow matching (FM) to train $g_{\theta}$. FM takes known distribution p and finds a path of probabilities to an unknown distribution q \@BBOPcitep\@BAP\@BBN(Lipman et al., 2023)\@BBCP. This probability path is discretized over $S\in\mathbb{N}$ steps, where S denotes the total number of discrete time intervals used to approximate the continuous flow from p to q. In our case, p is the Gaussian distribution and q is the distribution of the residual of the IPSL-DCPP. Training involves learning $\theta$ such that $g_{\theta}$ predicts $r_{t+\delta}$. The inputs to $g_{\theta}$ are the predicted state of the deterministic model $f_{\theta}(X_{t})$, the previous state $X_{t-\delta}$ and a residual noised according to a randomly chosen FM timestep $s$, $(1-s)r_{t+\delta}+s\epsilon$, where $\epsilon$ is noise sampled from a Gaussian distribution. During training, we sample s from a standard normal distribution, and use the sigmoid function as done in \@BBOPcite\@BAP\@BBNEsser et al. (2024)\@BBCP. See Figure 2 for a visualization of the process.
At each FM timestep $s$, the probability path is defined by a vector field that assigns a direction and magnitude to each point in our data to move between distributions. $g_{\theta}$ is updated to represent a vector field with the following loss function:
$\mathcal{L}=\mathbb{E}_{s\in\mathcal{U}(0,1),\epsilon\in\mathcal{N}(0,1)}\|(g_{\theta}(X_{t},f_{\theta}(X_{t}),X_{t-\delta},(1-s)r_{t+\delta}+s\epsilon)-(r_{t+\delta}-\epsilon)\|_{2}^{2}$ (2)
To recreate $X_{t+\delta}$ at a particular FM step $s$ we combine the output of the deterministic and generative model:
$X_{t+\delta,s+1}=f_{\theta}(X_{t}|X_{t-\delta})+g_{\theta}(X_{t},f_{\theta}(X_{t}),X_{t-\delta},\frac{(1-s)r_{t+\delta}+s\epsilon)}{\sigma})$ (3)
. We refer to \@BBOPcitep\@BAP\@BBN(Lipman et al., 2023; Esser et al., 2024)\@BBCP for more details on the flow matching process. There are two differences in the training between ArchesClimate and ArchesWeatherGen:
  • We do not do out-of-distribution finetuning (training on data outside of the deterministic training dataset) as we have a distribution shift over time that needs to be captured in ArchesClimate. Finetuning could overfit to a particular temporal period.
  • The FM target is a vector field instead of a full sample (see Section 2.3 for more details). Predicting the full sample instead of the vector field caused instability in generated states, and we therefore predict the vector field when doing FM.
IPSL-DCPP has been split as follows: every ten years, the 10-member ensembles initialized in the years 1989 and 1999 and 2009 are held out as validation, and ensembles initialized in the years 1969,1979,2010-2015 are held out as the test set. We chose this test set to best target the task of interpolation. See C for a comparison of different train/test splits. We train the deterministic model for 10 hours on 4 A100 GPUs, and we train the probabilistic model for 20 hours on 4 A100 GPUs.
2.4 Inference
Once both the deterministic and generative models are trained, ArchesClimate auto-regressively generates sequential states. To generate a state, the model needs to move from Gaussian noise $\epsilon$ to the data distribution of the following step. ArchesClimate takes $M\in\mathbb{N}$ FM steps during inference to go from $r_{t+\delta,0}$ to $r_{t+\delta,S}$ where $M$ is a hyperparameter set at inference time. At each FM inference step $m$, the model outputs $r_{t+\delta,\psi_{m+1}}$ with ArchesClimate until it reaches $r_{t+\delta,S}$ where $\psi$ consists of evenly spaced values from 0 to S with a step size of $\frac{S}{M}$.
$r_{t+\delta,\psi_{m+1}}=r_{t_{+}\delta,\psi_{m}}+(\psi_{m+1}-\psi_{m})g_{\theta}(r_{t_{+}\delta,\psi_{m}}|f_{\theta}(X_{t}),X_{t-\delta})$ (4)
Each step in the sampling process can be interpreted as taking a small step $(\psi_{m+1}-\psi_{m})$ in the direction given by the vector field generated with $g_{\theta}$, which is then added to the current state $r_{t_{+}\delta,\psi_{m}}$.
Once ArchesClimate has followed the probability path to the data distribution, it can take $r_{t+\delta}$ and use Equation 3 to build the full next state. This is then be used as input to generate subsequent states. See Figure 3 for a visualization of this process. To initialize ArchesClimate for inference, the model begins with states taken from IPSL-DCPP. To generate states for the years 1969-1979, we will initialize ArchesClimate with January 1969 and February 1969 from a random ensemble member of IPSL-DCPP.
Figure 2: Deterministic and Generative training schemes for ArchesClimate. It is necessary to have full trained $f_{\theta}$ before training $g_{\theta}$.
Figure 3: Sampling with ArchesClimate. Initial states and noise are given to $g_{\theta}$ and slowly shift from noise to the data distribution. The combined result of $f_{\theta}$ and $g_{\theta}$ are then used as input for the following timestep $t$.
3 Experiments
In this section, we describe the setup, evaluation methods and diagnostic tools used in experiments conducted with ArchesClimate. We then describe each experiment and its results.
3.1 Experimental Setup
To measure how well ArchesClimate captures the dynamics of IPSL-DCPP, we consider the following baseline. For each target ensemble, we select 5 members of the 10 members to serve as a baseline. We then compare both the remaining 5 ensemble members of IPSL-DCPP and a 5-member ArchesClimate ensemble to the baseline.
An alternative approach is to use climatology as a baseline, i.e. the monthly mean averaged over the training set; however, we found that climatology provided no useful signal. After approximately one year, climatology performed better than the held-out 5-member ensemble on variables with little to no climatic trend, but performed worse for variables with a discernible trend. As such, we opted for the more informative baseline described above.
For analyses requiring the anomaly of the predicted state, we calculate the anomaly by removing the monthly mean taken from the training period.
3.2 Evaluation Metrics and Diagnostic Tools
Continuous Ranked Probability Score (CRPS) is a metric used to evaluate the accuracy of probabilistic forecasts by measuring the difference between the predicted cumulative distribution function and the observed outcome. We use the same implementation of CRPS as in \@BBOPcite\@BAP\@BBNRasp et al. (2024)\@BBCP.
Rank Histograms are a diagnostic tool used to evaluate the consistency or calibration of an ensemble forecast by comparing it to another ensemble \@BBOPcitep\@BAP\@BBN(Hamill, 2001)\@BBCP. Using rank histograms, we evaluate whether an ensemble member generated with ArchesClimate can be interchanged with any member of the IPSL-DCPP ensemble. To calculate the rank of an ensemble member, we take the pixel-by-pixel values of the ensemble member and compare them to a target ensemble. We calculate the rank of the pixel value of the ensemble member compared to the rest of the target ensemble members at that pixel. We then take an average of the rank of each pixel across space and time. The histograms show the normalized frequency of the emulated ensemble member at that rank. We normalize the frequency by taking the mean and standard deviation of all frequencies across space and time. If the generated ensemble member can be interchanged with any member of an IPSL-DCPP ensemble, the histogram will be flat, showing an even distribution across all ranks.
Temporal Power Spectra (TPS) describes how the variance (or power) of a signal that changes with time is distributed among components that oscillate at specific rates (temporal frequencies, measured in cycles per unit time). This shows the strength of signals at varying timescales. We compute the TPS by calculating the temporal Fourier transform of the time series at each spatial location. We then take the positive frequencies and average spatially. We formalize this definition below:
[…]
$\mathrm{lat}=1,\dots,N_{y},\quad\mathrm{lon}=1,\dots,N_{x},\quad\mathcal{K}=\left\{\,k\;\middle|\;1\leq k\leq\left\lfloor\frac{T}{2}\right\rfloor=60\,\right\}.$
$X(k,\mathrm{lat},\mathrm{lon})=\sum_{n=0}^{119}x(n,\mathrm{lat},\mathrm{lon})\,e^{-i2\pi kn/120}.$ (5)
$\overline{\mathrm{PSD}}(k)=\frac{1}{N_{x}N_{y}}\sum_{lat=1}^{N_{y}}\sum_{lon=1}^{N_{x}}\left|X(k,lat,lon)\right|,\quad k\in\mathcal{K}$ (6)
Pearson Correlation Coefficient (PCC) measures the strength and direction of the linear relationship between two variables. It ranges from -1 to 1, where values close to 1 indicate a strong positive linear relationship, values close to -1 indicate a strong negative linear relationship, and values near 0 suggest little to no linear correlation between the variables.
Figure 4: Ensemble means of the full state (top) and Ensemble means for anomalies (bottom) of the Tropics (20° S – 20° N, 0° – 360° E) for the years 1969-1979. The dotted lines are the maximum and minimums for each ensemble mean, with the shaded area being +/- 1 standard deviation.
3.3 Regional Accuracy of ArchesClimate in the Tropics
By training on the full state of a given timestep, ArchesClimate can learn both the signal for the seasonal cycle and the anomaly, both of which are important aspects of decadal climate analysis \@BBOPcitep\@BAP\@BBN(Smith et al., 2019)\@BBCP. The seasonal cycle is the predictable annual pattern of changes caused by Earth’s position around the Sun, while anomalies are deviations from this expected seasonal average. Looking at anomalies is important in climate science because they highlight changes relative to a long-term average, making it easier to detect trends and patterns beyond natural variability. In this experiment, we qualitatively compare the performance of ArchesClimate to IPSL-DCPP in the Tropics (20°S–20°N, 0°–360°E) for both the full state and the anomaly of the full state. We use ArchesClimate to generate a 10-member ensemble and compare it to a 10-member ensemble of IPSL-DCPP for the test period initialized at 1969 for the Tropics. Each ensemble is 120 months (10 years), and we assess the stability and the regional accuracy of the generated states from ArchesClimate. We target the Tropics as they exhibit many important dynamics at climatic timescales \@BBOPcitep\@BAP\@BBN(Wang, 2019)\@BBCP. To show the performance in both ocean and atmosphere, we select a subset of variables emulated in ArchesClimate. Sea surface temperature (tos) and sea level pressure (psl) exemplify ocean dynamics, while total positive downward flux (net_flux) shows interactions between atmosphere and ocean. Air temperature at 700 hPa (ta) provides an example of atmospheric dynamics.
In Figure 4, ArchesClimate captures the seasonal cycle across the selected variables. When the seasonal cycle is removed, we can see that ArchesClimate produces anomalies similar to IPSL-DCPP. There are small deviations between ArchesClimate and IPSL-DCPP at the end of the 10 years in the temperature variables tos and ta, and significant variance around 1975 in IPSL-DCPP that is not present in ArchesClimate. As each ensemble member is auto-regressively generated, internal variability will be present in each ensemble member, providing deviations from IPSL-DCPP. ArchesClimate generated stable states for the 10-year period of the experiment and accurate anomalies of these stable states for the variables presented.
Figure 5: Variance (top) and CRPS (bottom) for different training schemes for the decade 1969-1979. residual-flow is described in Section 2.3. full-flow trains without any deterministic model. full-deterministic uses only the deterministic model to make predictions. The dotted line represents the 5-member IPSL-DCPP baseline.
3.4 Comparison of Training Schemes
ArchesClimate uses a combination of deterministic and generative models. To assess the training scheme outlined in Section 2.3 (hereby referred to as residual-flow), we compare it to several other schemes. First, we compare to a model where the residuals (as defined in Section 2.3) are trained with Denoising Diffusion Probabilistic Models (DDPM) instead of FM and call this residual-ddpm. DDPM gradually adds Gaussian noise to data, then learns to reverse this corruption to generate samples \@BBOPcitep\@BAP\@BBN(Ho et al., 2020)\@BBCP. We then make predictions using only a deterministic model and only a probabilistic model, named full-deterministic and full-flow, respectively. We compare these methods to our baseline outlined in Section 3.1.
As residual-ddpm is unable to generate stable states for longer than approximately 20 months, we do not include it in Figure 5. We see in Figure 5 that full-deterministic exhibits low variance compared to the other methods. Full-flow has similar variance to IPSL-DCPP but is less accurate than residual-flow across all variables. By combining the probabilistic and deterministic components (as shown in residual-flow), ArchesClimate makes much more accurate predictions and generates more variance than full-deterministic. We can see in tos that both full-deterministic and full-flow have high CRPS scores compared to IPSL-DCPP, but residual-flow has a similar CRPS score to IPSL-DCPP. The probabilistic model, when trained in tandem with the deterministic model, provides error correction and improves accuracy on top of increasing variance. We investigate methods to increase the variance of residual-flow in Section 3.10. Because FM learns a deterministic continuous-time flow, it requires fewer integration steps than DDPM at inference, making it more efficient and less prone to error accumulation. The implementation of FM is significantly simpler than DDPM, with fewer hyperparameters and implementation details and therefore helps FM perform better than DDPM.
3.5 CRPS and Variance of ArchesClimate
By computing the average CRPS and Variance over each period, we assess how the model responds to different forcings and different initializations. In this experiment, we quantify the performance of ArchesClimate using CRPS and Variance and compare ArchesClimate to our baseline averaged over three periods: 1969-1979, 1979-1989 and 2010-2020. We add vertically-averaged potential temperature at 0-2000m (thetaot2000) and several atmospheric variables (Omega (wap), geopotential (zg) and relative humidity (hur)) to the subset of variables already used in Section 3.3 to display a wider range of dynamics in ArchesClimate.
In Table 2, the CRPS for all variables except for thetaot2000 is lower in ArchesClimate than in IPSL-DCPP. The variance is consistently higher in IPSL-DCPP across all periods. As we predict the mean state deterministically and the residuals probabilistically, it is possible to increase the variance of the initial noise during the generation of the residual to increase the variance. We explore several ways to improve variance in Section 3.10. A lower CRPS score means that ArchesClimate is better at capturing the distribution of our baseline, the remaining 5 members of the IPSL-DCPP, than the first 5 members of the IPSL-DCPP. It is not clear that a lower CRPS score is better, as we want to replicate the dynamics of the IPSL-DCPP and therefore have similar CRPS to IPSL-DCPP. This ambiguity in CRPS necessitates more qualitative analysis of the generated states from ArchesClimate. Qualitative analysis is carried out in the subsequent experiments.
Variable CRPS Variance
ArchesClimate IPSL-DCPP ArchesClimate IPSL-DCPP
thetaot2000 (°C) 0.06 0.05 0.02 0.05
tos (°C) 0.27 0.27 0.26 0.51
psl (Pa) 87.98 106.39 30529.94 109017.79
net_flux (W/m-2) 7.25 7.97 270.42 544.49
zg 700 hPa (m) 8.38 10.03 242.58 938.19
wap 700 hPa (Pa/s-1) 1.07e-02 1.14e-02 6.53e-04 9.67e-04
ta 700 hPa (K) 0.41 0.46 0.68 1.58
hur 700 hPa (%) 2.78 3.02 32.40 55.38
Table 2: Comparison of CRPS and Variance between ArchesClimate and IPSL-DCPP. Scores are averaged over space and time for three periods: 1969-1979, 1979-1989 and 2010-2020. Both metrics keep the units shown in the table.
3.6 Interchangeability using Rank Histograms
Ensemble members generated by ArchesClimate should be able to be exchanged with members of IPSL-DCPP. We take a 10-year 10-member IPSL-DCPP ensemble initialized at 1969 for the North Atlantic (0°–65° N, 80°–0° W) and compare a 10-member ensemble of ArchesClimate one at a time for the same region and period. See Section 3.1 for a more detailed explanation of rank histograms. We chose the North Atlantic to complement the earlier assessment of the Tropics, as the North Atlantic exhibits important dynamics at the decadal timescale that differ from the Tropics \@BBOPcitep\@BAP\@BBN(Eade et al., 2014)\@BBCP. We use the same variable subset as in Section 3.5.
ArchesClimate produces relatively flat rank histograms for net_flux, wap, hur and ta, which indicate that the rank of the ArchesClimate member is likely to appear at each rank of the IPSL-DCPP ensemble. Bias in a rank histogram is represented by a skew to either end of the rank histogram. ArchesClimate often under-predicted values in the ocean variables tos and thetaot2000 and therefore was often ranked as the lowest. These histograms are good evidence that for several variables, our ensemble members are exchangeable with the IPSL-DCPP ensemble.
3.7 Comparison of Temporal Power Spectra
This experiment compares the Temporal Power Spectra (TPS) of both IPSL-DCPP and ArchesClimate to compare signals at different temporal frequencies. We look at the TPS for a 10-year 10-member ensemble initialized at 1969 for the North Atlantic. We use anomalies for all variables and use the same variable subset as in Section 3.3.
In Figure 6, there is a noticeable annual, seasonal and monthly signal for all variables. For all variables that have u-shaped rank histograms, ArchesClimate is underpowered in all but the strongest cycles. In psl (sea-level pressure), ArchesClimate correctly captures the strong annual cycle but under-represents everywhere else. This gives us insight into where ArchesClimate is unable to capture the variability of certain variables. It is clear in Figure 6 that the variables with lower variance are unable to generate enough power at frequencies outside of monthly, seasonal and annual cycles. This helps inform how to increase variance by including spectral information in the loss function. We investigate this idea in Section 3.10.
Figure 6: Diagnostics for the North Atlantic for years 1969-1979. On the left, Rank Histograms for ArchesClimate showing normalized frequency of the rank of a single ArchesClimate member in a 10-member ensemble of IPSL-DCPP. On the right, comparison of the temporal power spectral density of anomalies across time for ArchesClimate and IPSL-DCPP. The x-axis is logarithmic, with the smallest value being one month.
3.8 Calculating Linear Trend with Pearson Correlation Coefficient
We use Pearson Correlation Coefficient (PCC) to understand how well the ensemble mean of ArchesClimate captures spatial decadal trends of IPSL-DCPP. We calculate the per-pixel PCC for sea surface temperature anomalies, using a 10-year 5-member ensemble from ArchesClimate.
We show in Figure 7 that in regions with strong teleconnections, i.e. the North Atlantic, Pacific, and Indian Ocean, similar correlation over the decade as IPSL-DCPP. The persistence of these correlations in key basins indicates that ArchesClimate retains a similar level of decadal-scale predictive skill as IPSL-DCPP. ArchesClimate decorrelates throughout the decade in the high latitudes much more than IPSL-DCPP. In each of the three test periods, there is a decorrelation near Northern Canada and Greenland. ArchesClimate does not represent any sea-ice or ice dynamics, and is therefore limited in its ability to capture long-term dynamics of Arctic regions.
Figure 7: Pearson Correlation Coefficient (PCC) of tos (sea surface temperature) anomalies for several periods for both IPSL-DCPP and ArchesClimate. The PCC is calculated per pixel over the 10 years.
3.9 Spatial Anomaly Analysis in the North Atlantic
In previous experiments, we investigated the ability of ArchesClimate to capture regionally and globally averaged patterns (Section 3.3 and Section 3.5). We look again at the anomalies of sea surface temperature in the North Atlantic, but now to understand how well ArchesClimate captures the spatial signal of the seasonal cycle. We compute averages of seasonal sea surface temperature anomalies for the 10-year test period initialized at 1969 over the North Atlantic (0°–65° N, 80°–0° W). The seasonal means are MAM (March, April, May), JJA (June, July, August), SON (September, October, November) and DJF (December, January, February). We compare the results to the baseline outlined in Section 3.1.
Figure 8: Averages of seasonal anomalies of sea surface temperature over the North Atlantic (0°–65° N, 80°–0° W) for 1969-1979.
We can see that in Figure 8, ArchesClimate captures spatial patterns of the summer seasons. In the winter seasons, ArchesClimate produces colder states than IPSL-DCPP in the northern-most part of the Atlantic. This could be because the effect of the external forcings is much higher in the winter seasons; the mixed layer of the ocean is shallower in the summer months, relating to a weaker effect. In Figure 4, there is no clear indication that there is a bias in the winter seasons and that ArchesClimate can understand on aggregate seasonal patterns of variability.
3.10 Improving Variance in ArchesClimate
In this experiment, we explore ways to increase the variance lacking in ArchesClimate shown in Table 2. We compare three methods to improve variability. First, we multiply the initial noise at inference time by 1.1 (referred to as AC_noised) following ArchesWeatherGen \@BBOPcitep\@BAP\@BBN(Couairon et al., 2024)\@BBCP. The generative model takes the scaled noise distribution and translates this to a higher variance output distribution. Second, we apply noise scaling proportional to the difference in variance between the generated data and the target data for the validation period for each variable (AC_per_variable). Finally, we train another model with a spectral and image gradient loss (AC_updated_loss). To balance the weighting of the loss functions, we scale the gradient loss and spectral loss by 0.2. The loss function is as follows:
$\mathcal{L}_{\text{MSE}}=\mathbb{E}\left[(P_{i,j}-G_{i,j})^{2}\right]$ (7)
$\mathcal{L}_{\text{grad}}=\mathbb{E}\left[\left|(P_{i,j+1}-P_{i,j})-(G_{i,j+1}-G_{i,j})\right|^{2}\right]+\mathbb{E}\left[\left|(P_{i+1,j}-P_{i,j})-(G_{i+1,j}-G_{i,j})\right|^{2}\right]$ (8)
$\mathcal{L}_{\text{PSD}}=\mathbb{E}\left[\left(\log\left(|\mathcal{F}(P)|^{2}+\epsilon\right)-\log\left(|\mathcal{F}(G)|^{2}+\epsilon\right)\right)^{2}\right]$ (9)
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{MSE}}+0.2*\mathcal{L}_{\text{grad}}+0.2*\mathcal{L}_{\text{PSD}}$ (10)
Where:
$P$
is the prediction
$G$
is the ground truth
$P_{i,j}$
is the pixel at $i,j$
$\mathcal{F}$
denotes the 2D Fourier transform of the predicted image
$|\mathcal{F}|^{2}$
denotes the Power Spectra Density
In Figure 9 we can observe that both the noise scaling and per-noise scaling has little effect, and the alternate loss function successfully captures variance in tos and ta while improving significantly the variance in psl previously having low variance. The benefit here is twofold. The extra terms in the alternate loss function encourage ArchesClimate to pay attention more to spatial patterns and make the learning task much more difficult. While this new loss function improves variance in ArchesClimate, it reduces the accuracy of the model. The loss function increases CRPS for several variables, and balancing this tradeoff is left for further research.
Figure 9: Variance (top) and CRPS (bottom) of different techniques to increase variance in ArchesClimate. AC_noised increases initial noise at inference time by a factor of $1.$ AC_updated_loss uses an alternate loss function including gradient and spectral components. AC_per_variable scales the initial noise by the difference of variance in a generated ensemble versus the target ensemble for the test period for each variable.
4 Discussion and Conclusions
Our research aims to advance climate modeling by providing an inexpensive tool to generate ensemble members, enabling more robust probabilistic analysis. We propose ArchesClimate and demonstrate its use in generating ensemble members of the IPSL-CM6A-LR by training on climate model outputs and using forcings from input4mips to condition our model. We show that a combination of a deterministic and a generative machine learning model (based on ArchesWeatherGen) is an effective way to learn climate dynamics at low computational cost. We also find that the model can autoregressively produce climatically consistent 10-year sequences at a one-month timestep.
We evaluate ArchesClimate using both statistical and physical evaluations and find that it reliably reproduces key features of decadal climate variability. The model captures long-term correlations among ensemble members across regions that drive climate through teleconnections, while also reproducing seasonal spring and summer anomalies in close agreement with IPSL-DCPP. Across oceanic and atmospheric variables, ArchesClimate achieves comparable CRPS performance to IPSL-DCPP over multiple decades and replicates temporal power spectra consistent with major climate signals. Together, these results demonstrate that ArchesClimate is a reasonable approach for studying decadal climate variability.
While we train with data that is the output of a climate model, the model does not receive the initial conditions used to start each ensemble member of IPSL-DCPP. The impact of omitting the initialization state of the climate model is limited, as initialization noise has been shown to fade after a short period \@BBOPcitep\@BAP\@BBN(Smith et al., 2019)\@BBCP. In ArchesClimate, we create variation between ensemble members by generating states from different samples of Gaussian noise.
We include a holistic climate state in order to improve forecasting accuracy and generate more physically consistent states. In future work, however, possible that ArchesClimate could benefit from a more thorough set of forcings and input variables (e.g. sea ice, land-use). Similarly, a running average of the last decade’s climate could help the model as a sort of memory, especially for variables that operate on longer timescales \@BBOPcitep\@BAP\@BBN(Mignot et al., 2016)\@BBCP. By better constraining the model through these extra variables, we would expect to produce a more constrained output.
Our objective was to capture the underlying distribution of each physical variable at each timestep so that we can generate samples of the distribution of a given climate while still obeying dynamics that are resolved at shorter timescales. This task is difficult as the target distribution is hidden and the assessment of samples from the distribution requires expertise in climate science \@BBOPcitep\@BAP\@BBN(Mignot et al., 2016)\@BBCP. To aid the assessment of samples, we proposed a set of evaluations of both the statistics and the physical properties of the AI-generated ensembles that can be used in future evaluations of probabilistic climate model emulators. Further work can be done to find a base set of metrics and evaluations using already existing tools such as PCMDI, ESMValTool, climpred, and xMIP. With more thorough metrics, we can increase confidence in the ability of climate model emulators to augment climate models. Analyses can also be done to see if the climate model emulator adheres to conservation properties such as hydrostatic constraints and total water in atmospheric columns \@BBOPcitep\@BAP\@BBN(Sha et al., 2025; White et al., 2024; Watt-Meyer et al., 2024)\@BBCP.
Our work opens several new directions of research. Using members generated from ArchesClimate, we believe it will be possible to temporally and spatially downscale using similar generative methods to enable analysis at higher resolutions. Further work also includes training on longer experiments such as the Coupled Model Inter-comparison Project experiment historical simulations that span several hundred years. By extending ArchesClimate to multi-decadal climate simulations, we can assess if it can emulate inter-annual variability at long timescales \@BBOPcitep\@BAP\@BBN(Jain et al., 2023)\@BBCP. Expanding the forcings already used in ArchesClimate, interpolation between longer experiments can be investigated to help climate scientists explore previously untested future climate scenarios.
Another possibility is to explore recent advances in generative methods, for which inference time can be done in a fraction of the time, sometimes using only a single inference step \@BBOPcitep\@BAP\@BBN(Hess et al., 2025; Schmitt et al., 2024)\@BBCP, which would cut the cost of generating samples by an order of magnitude. Our results suggest that emulating climate states at a monthly resolution is an efficient way to predict long-term climate dynamics. Further work is needed to explore the limits of jointly predicting multiple dynamics at monthly time scales. It remains an open question whether ArchesClimate is primarily capturing the underlying processes that operate at shorter timescales, or whether some of its skill at monthly prediction may arise from correlations present at those scales.
ArchesClimate offers a powerful and efficient complement to traditional climate modeling approaches. By leveraging machine learning, ArchesClimate produces decadal climate predictions with comparable accuracy at a fraction of the computational cost of running a climate model. This advancement enables broader access to probabilistic climate projections and supports more timely, informed decision-making in response to climate change.
Appendix A Derivation of net_flux
This is a derivation of the variable net_flux that is used in ArchesClimate. We define net_flux as:
net_flux=rsusrsds+rlusrlds+hfss+hflsnet\_flux=rsus-rsds+rlus-rlds+hfss+hfls (11)
rsus
Surface Upwelling Shortwave Radiation
rsds
Surface Downwelling Shortwave Radiation
rlus
Surface Upwelling Longwave Radiation
rlds
Surface Downwelling Longwave Radiation
hfss
Surface Upward Sensible Heat Flux
hfls
Surface Upward Latent Heat Flux
Appendix B Long-term Forcing Response
We generate states for 50 years to test the ability of ArchesClimate to respond to external forcings and remain stable over 50 years. We compare these states to generated states that are conditioned with forcings for the year 1969 are repeated every year for 50 years. We take the first ensemble of the validation period (the ensemble initialized at 1969) and the first ensemble of the test period (the ensemble initialized at 2010) to mark the beginning and the end of the 50-year rollout.
Figure 10 shows the results for sea surface temperature, where ArchesClimate is much closer to the trend of the dataset than the rollout with repeated forcings. Even with limited forcings, there is a noticeable effect on sea surface temperature. By expanding the scope of the forcings, we may be able to capture more complex long-term dynamics.
Figure 10: Yearly means of a 50-year rollout of ArchesClimate with and without constant forcings for tos (sea surface temperature). Two IPSL-DCPP decades are used for reference.
Appendix C Alternative Train/Test Splits
In the train/test split of IPSL-DCPP described in Section 2.3, there is temporal overlap in the training and test sets. Here, we look at a train/test split that leaves a test set temporally apart from the train set. To do this, we use the initialization years 1960-2000, so the last year of training is 2009, which we refer to in Figure 11 as alternative. We can then compare a model that has seen the years 2010-2020 to a model that has not. We generate a 10-member ensemble initialized in 2010 for both models.
We compare CRPS in Figure 11. There is a noticeable difference in CRPS between ArchesClimate and IPSL-DCPP for the variables tos and ta towards the end of the decade. ArchesClimate performs much better when it has seen samples from the period of generation. Further investigation is needed to understand if a more comprehensive set of external forcings will improve the ability of ArchesClimate to extrapolate to unseen futures. We use the original train/test split for several reasons: our goal in this research is to augment the IPSL-CM6A-LR and not to extend the dataset beyond its current period. As well, the current train/test split allows for temporal comparison throughout the dataset, providing a more thorough analysis.
Figure 11: Comparison of CRPS for different train/test splits. “original” indicates the training scheme outlined in Section 2.3 and “alternative” indicates the training scheme where training and test contain no temporal overlap. The dotted line is the 5-member IPSL-DCPP baseline.
Open Research Section
Code for the project can be found at https://github.com/INRIA/geoarches. Training data can be found at https://esgf-node.ipsl.upmc.fr/projects/cmip6-ipsl/.
Acknowledgements.
G. Clyne, G. Couairon, and C. Monteleoni were primarily supported by the Choose France Chair in AI, from the French government. The authors thank Juliette Mignot, David Landry, Clément Dauvilliers, and Renu Singh for several discussions about the project. To process the data from IPSL, this study benefited from the IPSL mesocenter ESPRI facility, which is supported by CNRS, UPMC, Labex L-IPSL, CNES and Ecole Polytechnique.